# Estimation of discrete distributions in relative entropy, and the deviations of the missing mass

### v1

@@ Abstract
We study the problem of estimating a distribution over a finite alphabet from an i.i.d. sample, with accuracy measured in relative entropy (Kullback-Leibler divergence). While optimal expected risk bounds are known, high-probability guarantees remain less well-understood. First, we analyze the classical Laplace (add-$1$) estimator, obtaining matching upper and lower bounds on its performance and showing its optimality among confidence-independent estimators. We then characterize the minimax-optimal high-probability risk achievable by any estimator, which is attained via a simple confidence-dependent smoothing technique. Interestingly, the optimal non-asymptotic risk contains an additional logarithmic factor over the ideal asymptotic risk. Next, motivated by scenarios where the alphabet exceeds the sample size, we investigate methods that adapt to the sparsity of the distribution at hand. We introduce an estimator using data-dependent smoothing, for which we establish a high-probability risk bound depending on two effective sparsity parameters. As part of the analysis, we also derive a sharp high-probability upper bound on the missing mass.
@@ 1.1 Problem setting
An important feature of the Kullback-Leibler divergence, compared to other common divergences between probability distributions such as the total variation and Hellinger distances, is that it penalizes significant underestimation of true frequencies. To consider an extreme case, if the estimator $\widehat{P}$ assigns a probability $\widehat{p}_{j}=0$ to a class $j$ with probability $p_{j}\neq 0$, then the Kullback-Leibler divergence $\mathrm{KL}(P,\widehat{P}_{n})$ is infinite. This aligns with the needs of various applications: for instance, in a forecasting context where classes correspond to different outcomes, assigning a probability of 00 to outcomes that are in fact possible would constitute a severe underestimation of the underlying uncertainty. In addition, in the context of language models, one may be interested in generating new sentences that are not present in the training corpus; this requires assigning positive probabilities to sequence of words that have not been observed.
@@ 1.2 Existing guarantees
Finally, the best available high-probability guarantees on the Laplace estimator are due to Canonne, Sun and Suresh [CSS23], who showed that for some absolute constant $c>0$, with probability at least $1-\delta$,
@@
(where we used (5) and that $\sqrt{d}\log^{5/2}d\lesssim d$). In particular, this removes a $\sqrt{\log(1/\delta)}$ factor in the deviation term compared to (9), leading to optimal guarantees in the larger range $\log(1/\delta)\lesssim d^{1/5}$. Nevertheless, the bounds still deteriorates in the regime $\log(1/\delta)\gg d^{1/5}$.
@@
In many regimes of interest, this constitutes the best known high-probability guarantee in the literature, for any estimator. On the other hand, this result raises important questions. First, the bound (11) features an additional $\log n$ factor compared to the asymptotic rate (3), which can be avoided in some regimes (e.g., in light of (10)), leaving open the question of the best possible statistical guarantees. Second, the estimator $\widehat{P}_{n}^{\mathrm{HZC}}$ is computationally expensive: it requires integrating non-log-concave functions over the probability simplex, the cost of which a priori scales exponentially in the dimension $d$ via a grid-based approach. This raises the question of whether or not the problem of high-probability estimation of discrete distributions exhibits a statistical-computational trade-off, that is, if a computational cost exponential in $d$ is necessary to achieve a guarantee as strong as (11).
@@ 1.4 Related work
As part of our analysis, we study the tail behavior of the “missing mass”. We refer to this Section 5 for a discussion of existing high-probability bounds on this quantity.
@@ 2.1 Upper bound for the Laplace estimator
We refer to Section 6 for a description of the main tools in the proof of Theorem 1, which also serve for high-probability upper bounds stated in subsequent sections. Roughly speaking, the key step in the analysis is to control the contribution to the error $\mathrm{KL}(P,\widehat{P}_{n})$ of classes $j=1,\dots,d$ for which the Laplace estimate significantly underestimates the true probability, as such classes may lead to a potentially large errors.
@@ 2.2 Lower bound for confidence-independent estimators
The proof of Theorem 2 (based on Lemma 12) can be found in Section ref. The main idea of the proof is that, in order to achieve the guarantee (17), the estimator $\widehat{P}_{n}$ cannot be “too far off” the empirical distribution $\overline{P}_{n}$. But in this case, it may significantly underestimate the probability of some classes with small probability $\delta$, leading to the lower bound (18).
@@ 4.2.2 Upper bound in expectation for sparse distributions
The estimator $\widehat{P}_{n}^{\mathrm{ad}}$ defined by (31) can be viewed as an adaptive modification of the Laplace estimator (4), where the regularization parameter $\widehat{\lambda}$ is chosen in a data-dependent manner, so as to adapt to the “shape” of the distribution $P$. In the case where $n\gtrsim d$ and the distribution is “dense”, namely puts significant probability $p_{j}\gtrsim 1/d$ to all classes, then typically $D_{n}$ is of order $d$ and thus $\widehat{\lambda}\asymp 1$, hence the estimator behaves similarly to the Laplace estimator. On the other hand, when the distribution is highly sparse, then typically $D_{n}\ll d$ and thus the regularization parameter $\widehat{\lambda}\ll 1$ is much smaller than for the Laplace estimator—it may be as low as $1/d$, in the extreme case where only one class appears in the sample. The specific choice of tuning parameter $\widehat{\lambda}=D_{n}/d$ is motivated by the decomposition of the risk in Lemma 3 below.
@@ 4.2.3 High-probability upper bounds for sparse distributions
$n\gtrsim N_{\mathsf{max}}(d,\varepsilon,\delta)=\frac{d+\log(d)\log(1/\delta)}{\varepsilon}$
@@ 5 High-probability bound on the missing mass
Given a distribution $P$ and an i.i.d. sample $X_{1},\dots,X_{n}$ from $P$, we define the missing mass $M_{n}=\sum_{j=1}^{d}p_{j}\bm{1}(N_{j}=0)$ as the total mass under $P$ of classes that do not appear in the sample.
@@
Before stating our main result, we first clarify that the behavior of the expected missing mass is essentially governed by the behavior of the parameter $s_{n}^{\circ}(P)$ (Definition 2), a fact we alluded to in Section 4.
@@
(Specifically, from (41) we obtain an additive term in $3e^{1-0.15n}\leqslant 10e\log(1/\delta)/n$.) In addition, as will be discussed below, the deviation term in $\log(1/\delta)/n$ is necessary, except in situations where the probabilities in $P$ exhibit a significant gap; and in this case, the bound of Theorem 6 admits a simple strengthening that addresses its sub-optimality.
@@
    $n\geqslant\frac{1}{2}\max\bigg{\{}N_{\mathsf{exp}}(\varepsilon),\frac{\log(1/\delta)}{\overline{\varepsilon}(P,\varepsilon)}\bigg{\}}\,.$
@@ 6.2 Upper bound in Hellinger distance
We note in passing that this bound is of the right order of magnitude, as one may check that $\mathbb{E}_{P}[\sum_{j=1}^{d}(\sqrt{\overline{p}_{j}}-\sqrt{p_{j}})^{2}]\asymp s_{n}(P)/n$.
@@ 7.1 Proof of Theorem 2
$\mathbb{P}_{P}\bigg{(}\mathrm{KL}(P,\widehat{P})\geqslant\frac{\log(1/\delta)\log\log(1/\delta)}{10n}\bigg{)}\geqslant\delta\,.$ (90)
@@
In addition, under $E$ one has $\widehat{P}=Q$, thus

### v2

@@ Abstract
We study the problem of estimating a distribution over a finite alphabet from an i.i.d. sample, with accuracy measured in relative entropy (Kullback-Leibler divergence). While optimal expected risk bounds are known, high-probability guarantees remain less well-understood. First, we analyze the classical Laplace (add-one) estimator, obtaining matching upper and lower bounds on its performance and showing its optimality among confidence-independent estimators. We then characterize the minimax-optimal high-probability risk, which is attained via a simple confidence-dependent smoothing technique. Interestingly, the optimal non-asymptotic risk exhibits an additional logarithmic factor over the ideal asymptotic risk. Next, motivated by scenarios where the alphabet exceeds the sample size, we investigate methods that adapt to the sparsity of the distribution at hand. We introduce an estimator using data-dependent smoothing, for which we establish a high-probability risk bound depending on two effective sparsity parameters. As part of the analysis, we also derive a sharp high-probability upper bound on the missing mass.
@@ 1.1 Problem setting
An important feature of the Kullback-Leibler divergence, compared to other common divergences between probability distributions such as the total variation and Hellinger distances, is that it penalizes significant underestimation of true frequencies. To consider an extreme case, if the estimator $\widehat{P}_{n}$ assigns a probability $\widehat{p}_{j}=0$ to a class $j$ with probability $p_{j}\neq 0$, then the Kullback-Leibler divergence $\mathrm{KL}(P,\widehat{P}_{n})$ is infinite. This aligns with the needs of various applications: for instance, in a forecasting context where classes correspond to different outcomes, assigning a probability of 00 to outcomes that are in fact possible would constitute a severe underestimation of the underlying uncertainty. In addition, in the context of language models, one may be interested in generating new sentences that are not present in the training corpus; this requires assigning positive probabilities to sequence of words that have not been observed.
@@ 1.2 Existing guarantees
Finally, the best available high-probability guarantee on the Laplace estimator is due to Canonne, Sun and Suresh [CSS23], who showed that for some absolute constant $c>0$, with probability at least $1-\delta$,
@@
(where we used (5) and that $\sqrt{d}\log^{5/2}d\lesssim d$). In particular, this removes a $\sqrt{\log(1/\delta)}$ factor in the deviation term compared to (9), leading to optimal guarantees in the larger range $\log(1/\delta)\lesssim d^{1/5}$. Nevertheless, the bound still deteriorates in the regime $\log(1/\delta)\gg d^{1/5}$.
@@
In many regimes of interest, this constitutes the best known high-probability guarantee in the literature, for any estimator. On the other hand, this result raises important questions. First, the bound (11) features an additional $\log n$ factor compared to the asymptotic rate (3), which can be avoided in some regimes (e.g., in light of (10)), leaving open the question of the best possible statistical guarantees. Second, the estimator $\widehat{P}_{n}^{\mathrm{HZC}}$ is computationally involved: it requires integrating certain functions over the probability simplex, the cost of which appears to be super-linear in $\max(d,n)$ via a sampling approach. This raises the question of whether or not the problem of high-probability estimation of discrete distributions exhibits a statistical-computational trade-off, that is, if a computational cost super-linear in $n$ is necessary to achieve a guarantee as strong as (11).
@@ 1.4 Related work
As part of our analysis, we study the tail behavior of the “missing mass”. We refer to Section 5 for a discussion of existing high-probability bounds on this quantity.
@@ 2.1 Upper bound for the Laplace estimator
We refer to Section 6 for a description of the main tools in the proof of Theorem 1, which also serve for high-probability upper bounds stated in subsequent sections. Roughly speaking, the key step in the analysis is to control the contribution to the error $\mathrm{KL}(P,\widehat{P}_{n})$ of classes $j=1,\dots,d$ for which the Laplace estimate significantly underestimates the true probability, as such classes may lead to large errors.
@@ 2.2 Lower bound for confidence-independent estimators
The proof of Theorem 2 (based on Lemma 12) can be found in Section 7.1. The main idea of the proof is that, in order to achieve the guarantee (17), the estimator $\widehat{P}_{n}$ cannot be “too far” from the empirical distribution $\overline{P}_{n}$. But in this case, it may significantly underestimate the probability of some classes with small probability $\delta$, leading to the lower bound (18).
@@ 4.2.2 Upper bound in expectation for sparse distributions
The estimator $\widehat{P}_{n}^{\mathrm{ad}}$ defined by (31) can be viewed as an adaptive modification of the Laplace estimator (4), where the regularization parameter $\widehat{\lambda}$ is chosen in a data-dependent manner, so as to adapt to the “shape” of the distribution $P$. In the case where $n\gtrsim d$ and the distribution is “dense”, namely puts significant probability $p_{j}\gtrsim 1/d$ to all classes, then typically $D_{n}$ is of order $d$ and thus $\widehat{\lambda}\asymp 1$, hence the estimator behaves similarly to the Laplace estimator. On the other hand, when the distribution is highly sparse, then typically $D_{n}\ll d$ and thus the regularization parameter $\widehat{\lambda}\ll 1$ is much smaller than for the Laplace estimator—it may be as low as $1/d$, in the extreme case where only one class appears in the sample. The specific choice of tuning parameter $\widehat{\lambda}=D_{n}/d$ is motivated by the risk decomposition in Lemma 3 below.
@@ 4.2.3 High-probability upper bounds for sparse distributions
$n\gtrsim\frac{d+\log(d)\log(1/\delta)}{\varepsilon}$
@@ 5 High-probability bound on the missing mass
Given a distribution $P$ and an i.i.d. sample $X_{1},\dots,X_{n}$ from $P$, The missing mass $M_{n}=\sum_{j=1}^{d}p_{j}\bm{1}(N_{j}=0)$ is the total mass under $P$ of classes that do not appear in the sample.
@@
Before stating our main result, we first clarify that the behavior of the expected missing mass is essentially governed by the parameter $s_{n}^{\circ}(P)$ (Definition 2), a fact we alluded to in Section 4.
@@
(Specifically, from (41) we obtain an additive term in $3e^{1-0.15n}\leqslant 20e\log(1/\delta)/n$.) In addition, as will be discussed below, the deviation term in $\log(1/\delta)/n$ is necessary, except in situations where the probabilities in $P$ exhibit a significant gap; and in this case, the bound of Theorem 6 admits a simple strengthening that addresses its sub-optimality.
@@
    $n\geqslant\frac{1}{2}\max\bigg{\{}N_{\mathsf{exp}}(P,\varepsilon),\frac{\log(1/\delta)}{\overline{\varepsilon}(P,\varepsilon)}\bigg{\}}\,.$
@@ 6.2 Upper bound in Hellinger distance
We note in passing that this bound is of the right order of magnitude, as one may check that $\mathbb{E}_{P}[\sum_{j=1}^{d}(\sqrt{\overline{p}_{j}}-\sqrt{p_{j}})^{2}]\asymp s_{n}(P)/n$ whenever $\max_{j}p_{j}$ is bounded away from $1$.
@@ 7.1 Proof of Theorem 2
$\mathbb{P}_{P}\bigg{(}\mathrm{KL}(P,\widehat{P}_{n})\geqslant\frac{\log(1/\delta)\log\log(1/\delta)}{10n}\bigg{)}\geqslant\delta\,.$ (90)
@@
In addition, under $E$ one has $\widehat{P}_{n}=Q$, thus